If one-half a number is subtracted from five-sixths of the number, the difference is 6. Find the number.
Thank you.

Answer by stanbon(75887)

Let the number be "x".
EQUATION:
(5/6)x - (1/2)x = 6
x[(5/6)-(3/6)]=6
x[1/3]=6
x=18
Cheers,
Stan H.